OpenStudy (anonymous):

the model for an object falling is given by v(t) = (32.2/k) - (32.2/k)(e^-kt) solve k so it takes 2.6 seconds for the object to hit the ground

OpenStudy (anonymous):

The formula is \[v(t)=\frac{32.3}{k}-\frac{32.3}{k}(e^{-kt})\] Correct?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

Hehe cool. So then our \(t\)-value is 2.6 and assuming \(v(t)\) Is the height of the object from the ground, then we have \(v(t)=0\). Is \(v(t)\) the height of the object from the ground?

OpenStudy (anonymous):

It says that it falls from 100 feet so I'm assuming so.

OpenStudy (anonymous):

Okay. So we can make the formula: \[0=\frac{32.3}{k}-\frac{32.3}{k}(e^{-2.6k})\] And we must isolate for \(k\).

OpenStudy (anonymous):

Wouldn't v(t) be 2.6 since we are substituting 2.6 for t?

OpenStudy (anonymous):

No...because \(t=2.6\) does not mean \(v(t)=2.6\)

OpenStudy (anonymous):

Ok understood

OpenStudy (anonymous):

I get 13.46 for K

OpenStudy (anonymous):

So...we can simplify a little bit. \[0=\frac{32.3}{k}(1-e^{-2.6k})\] Since we know that \(\frac{32.3}{k}\) Will never be zero, then, through the property of zero, we know that \[1-e^{-2.6k}=0\] And we can isolate for k here \[e^{-2.6k}=1\] \[-2.6k=ln(1)\]...Ermmm wait no something is wrong. I believe that I substituted something in the wrong place...Haha my bad. Something is wrong with this...ill end up with k=0 and that's wrong. I feel like there is something more to this question. I wish I could figure it out, but I must go. Ill look at this tomorrow at a similar time.